Wednesday, March 11, 2020, 2:00 pm EDT

From Hall A Wiki
Jump to: navigation, search

Back to Main Compton Page >> Compton Meetings

*********BLUEJEANS CALL-IN INSTRUCTIONS*********
----------------------------------------------------
1. Dial Toll-Free Number: 1-888-240-2560 (U.S. & Canada)
2. International participants dial:
Toll Number: 1-408-740-7256
3. Enter meeting ID:140287970

Desktop sharing:
https://bluejeans.com/140287970 

Agenda

Meeting room = Counting House Meeting Room

General

  • Systematic checks - scheduling

Electron detector

  • QWAD installation/issues

Laser

  • Laser polarization systematic tests (updated) [1],[2]

Photon Detector

  • Spectrum fitting for collimator offset (low threshold, high threshold)
  • Strange feature in recent full gain runs
  • LED run

Analysis

  • Online analysis results
  • Polarization vs. time


Minutes

Mark, Dave, Alex, Cip, Iris, JC, Kent, Don, Jinlong,

  1. General
    • Dave shows polarization versus time for Hall C. Nice!
      • There are a number of inconsistencies, Hall C Moller measurements requires a different launch angle and linac imbalance to agree.
  2. Laser Systematic checks
    • Dave will look at a 10 degree square in HWP-QWP to find a setting to use that doesn't have too big a RRPD signal.
  3. Photon detector
    1. Issues with low energy part of the spectrum. Is this due to changes in discriminator-trigger path. Can we mitigate this by using the photon signal in the electron detector DAQ.
    2. LED run didn't work. Various things can be tried.
    3. Jaws wide open position seems to clip a bit at the top (about the same amount as the nominal clips at the bottom.)
    4. Finger scans.
      1. Inconclusive at the moment. New scans will be done.